Wesco Incorporated’s only product is a combination fertilizer/weedkiller called GrowNWeed. GrowNWeed is sold nationwide to retail nurseries and garden stores.

Zwinger Nursery plans to sell a similar fertilizer/weedkiller compound through its regional nursery chain under its own private label. Zwinger does not have manufacturing facilities of its own, so it has asked Wesco (and several other companies) to submit a bid for manufacturing and delivering a 26,000-pound order of the private brand compound to Zwinger. While the chemical composition of the Zwinger compound differs from that of GrowNWeed, the manufacturing processes are very similar.

The Zwinger compound would be produced in 1,000-pound lots. Each lot would require 37 direct labor-hours and the following chemicals:

Chemicals Quantity in Pounds
AG-5 320
KL-2 290
CW-7 190
DF-6 200

The first three chemicals (AG-5, KL-2, and CW-7) are all used in the production of GrowNWeed. DF-6 was used in another compound that Wesco discontinued several months ago. The supply of DF-6 that Wesco had on hand when the other compound was discontinued was not discarded. Wesco could sell its supply of DF-6 at the prevailing market price less $0.08 per pound selling and handling expenses.

Wesco also has on hand a chemical called BH-3, which was manufactured for use in another product that is no longer produced. BH-3, which cannot be used in GrowNWeed, can be substituted for AG-5 on a one-for-one basis without affecting the quality of the Zwinger compound. The BH-3 in inventory has a salvage value of $470.

Inventory and cost data for the chemicals that can be used to produce the Zwinger compound are shown below:

Raw Material Pounds in
Inventory
Actual Price
per Pound
When
Purchased
Current
Market
Price per
Pound
AG-5 24,000 $ 0.61 $ 0.71
KL-2 4,800 $ 0.45 $ 0.50
CW-7 8,700 $ 1.23 $ 1.43
DF-6 3,700 $ 0.50 $ 0.55
BH-3 5,400 $ 0.60 (Salvage)

The current direct labor wage rate is $18 per hour. The predetermined overhead rate is based on direct labor-hours (DLH). The predetermined overhead rate for the current year, based on a two-shift capacity with no overtime, is as follows:

Variable manufacturing overhead $ 4.60 per DLH
Fixed manufacturing overhead 7.70 per DLH
Combined predetermined overhead rate $ 12.30 per DLH

Wesco’s production manager reports that the present equipment and facilities are adequate to manufacture the Zwinger compound. Therefore, the order would have no effect on total fixed manufacturing overhead costs. However, Wesco is within 140 hours of its two-shift capacity this month. Any additional hours beyond the 140 hours must be done in overtime. If need be, the Zwinger compound could be produced on regular time by shifting a portion of GrowNWeed production to overtime. Wesco’s direct labor wage rate for overtime is $27 per hour. There is no allowance for any overtime premium in the predetermined overhead rate.

Required:

1. Wesco has decided to submit a bid for the 26,000 pound order of Zwinger’s new compound. The order must be delivered by the end of the current month. Zwinger has indicated that this is a one-time order that will not be repeated. Calculate the lowest price that Wesco could bid for the order and still exactly cover its incremental manufacturing costs.

2. Refer to the original data. Assume that Zwinger Nursery plans to place regular orders for 26,000-pound lots of the new compound. Wesco expects the demand for GrowNWeed to remain strong. Therefore, the recurring orders from Zwinger would put Wesco over its two-shift capacity. However, production could be scheduled so that 60% of each Zwinger order could be completed during regular hours. As another option, some GrowNWeed production could be shifted temporarily to overtime so that the Zwinger orders could be produced on regular time. Current market prices are the best available estimates of future market prices.

Wesco’s standard markup policy for new products is 40% of the full manufacturing cost, including fixed manufacturing overhead. Calculate the price that Wesco, Inc., would quote Zwinger Nursery for each 26,000 pound lot of the new compound, assuming that it is to be treated as a new product and this pricing policy is followed.
Answer #1

Answer for 1:

a. Calculation of Material Cost:

Given,

Raw Material Pounds in
Inventory
Actual Price
per Pound
When
Purchased
Current
Market
Price per
Pound
AG-5 24,000 $ 0.61 $ 0.71
KL-2 4,800 $ 0.45 $ 0.50
CW-7 8,700 $ 1.23 $ 1.43
DF-6 3,700 $ 0.50 $ 0.55
BH-3 5,400 $ 0.60 (Salvage)
Component Quantity in Pounds Relevant Cost Amount ($) Amount ($)
BH-3 5400 Salvage Value 470
AG-5 2920 Purchased Value =2920*0.61 1781.2
KL-2 4800 Purchased Value =4800*0.45 2160
KL-2 2740 Market Value =2740*0.50 1370
CW-7 4940 Purchased Value =4940*1.23 6076.2
DF-6 3700 Net Realisable Value =3700*(0.55-0.08) 1739
DF-6 1500 Market Value =1500*0.55 825
Total 14421.4

The relevant cost for material requirement which can be served through the existing inventory is taken at the Purchase Cost and relevant cost for the remaining quantity is taken at the current market prices.

In case of the BH-3 Component which can be substituted with the AG-5 on one-to-one is taken at the salvage value being the relevant cost.

b. Calculation of the Labor Cost

The total labor hours for the production of the 26000 pound order is 962 DLH (37 per lot * 26). But the regular labor hours available are 140 Hours and hence the remaining hours are to be done on over-time.

Labor Cost = 140*18 + 822*27 = 2520 + 22194 = $ 24714.

c. Calculation of Variable Overheads

Variable overheads = 962*4.60 = $ 4425.2

d. As it is assumed that this a one time order the fixed manufacturing overheads need not be charged as this order does not require any incremental fixed costs and is well within the capacity.

Final Answer: The lowest price that Wesco could bid for the order and still exactly cover its incremental manufacturing costs is the total of a, b and c i.e., $ 14421.4 + $ 24714 + $ 4425.2 = $ 43560.6

Answer for 2:

a. Calculation of Material Cost:

As the it is assumed that the order of 26,000 pounds comes on a regular basis the relevant cost of materials will the the future purchase cost and is given that current market prices are the best available estimates of future market prices. So the current market prices of the components will be the relevant cost.

Component Quantity in Pounds Amount ($) Amount ($)
AG-5 8320 =8320*0.71 5907.2
KL-2 7540 =7540*0.50 3770
CW-7 4940 =4940*1.43 7064.2
DF-6 5200 =5200*0.55 2860
Total 19601.4

b. Calculation of the Labor Cost

The total labor hours for the production of the 26000 pound order is 962 DLH (37 per lot * 26). As it is assumed that 60% of the regular labor hours will be used for the Wesco Inc., order, the remaining 40% DLH will be calculated on over-time basis.

Labor Cost = 577*18 + 385*27 = 10386 + 10395 = $ 20781.

c. Calculation of Manufacturing Overhead

As the it is assumed that the order of 26,000 pounds comes on a regular basis, the entire manufacturing overheads, both variable and fixed will be charged.

Manufacturing overheads = 962*12.3 = $ 11832.6

d. Calculation of the Mark-up on full manufacturing cost:

As Wesco’s standard markup policy for new products is 40% of the full manufacturing cost, including fixed manufacturing overhead, the Mark-up is 40% on (a+b+c) i.e., 40% of (19601.4+20781+11832.6) = 40% of $ 52,215 = $ 20,886

Final Answer:

The price that Wesco, Inc., would quote Zwinger Nursery for each 26,000 pound lot of the new compound, assuming that it is to be treated as a new product and this pricing policy is followed, is $ 52,215 + $ 20,886 = $ 73.101